OpenStudy (anonymous):

A wholesaler buys plastic lamps for $18 each and sells them for $30.60.what is the percent markup based on cost?

OpenStudy (***[isuru]***):

markup value = selling price - buying price = 30.6 - 18 \[percentage \ markup = \frac{ markup \ value }{ cost \ price} \times 100\] can u do it now ?

OpenStudy (***[isuru]***):

cost price = buying price \[percent \ markup = \frac{ 30.6 -18 }{ 18 } \times 100 = \frac{ 12.6 }{ 18 } \times 100 = ?\]

OpenStudy (anonymous):

thank u

OpenStudy (***[isuru]***):

u r welcome :)